# Universal projection theorems with applications to multifractal analysis and the dimension of every ergodic measure on self-conformal sets simultaneously

### v2

@@ 7.2. Gibbs measures
[…]
@@
$[\omega|_{n+1}]\subset B(\omega,r)\subset[\omega|_{n}]\text{ and }B(\omega,2r)\subset[\omega|_{n-N}].$
@@ 8. Multifractal spectrum of self-similar measures
Finally, we prove Theorem 3.14 as an application of the above results. It relies on two lemmas
@@
Let ${\mathcal{F}}^{\lambda},\lambda\in U$ be a continuous family of $C^{1+\theta}$ conformal IFS on $V\subset\mathbb{R}^{d}$. Assume that $\eta$ is a measure on $U$ such that the transversality condition (A3) is satisfied (for the family of corresponding metrics $\rho_{\lambda}$ as defined in (3.3)). Then for $\eta$-a.e. $\lambda\in U$ such that $s({\mathcal{F}}_{\lambda})<d$ the following holds: for every $\mu,\nu\in{\mathcal{E}}_{\sigma}(\Sigma)$ such that $\nu$ is a $k$-step Markov measure
(8.1) $d(\Pi_{\lambda}\nu,x)=\frac{h(\mu)-h(\mu\|\nu)}{\chi(\mu,\lambda)}\text{ for $\Pi_{\lambda}\mu$-almost every $x$.}$
@@
It is clear that for every $\lambda\in U$
$\overline{d}(\Pi_{\lambda}\nu,x)=\limsup_{n\to\infty}\frac{\log\Pi_{\lambda}\nu\left(B(\Pi_{\lambda}(u),\|(f^{\lambda}_{u|_{n}})^{\prime}\|)\right)}{\log\|(f^{\lambda}_{u|_{n}})^{\prime}\|}\leq\limsup_{n\to\infty}\frac{\log\nu\left([u|_{n}]\right)}{\log\|(f^{\lambda}_{u|_{n}})^{\prime}\|}=\frac{h(\mu)-h(\mu\|\nu)}{\chi(\mu,\lambda)},$
where $f^{\lambda}_{u|_{n}}=f^{\lambda}_{u_{1}}\circ\cdots\circ f^{\lambda}_{u_{n}}$, and so the upper bound in (8.1) follows by Theorem 7.5, the Shannon-McMillan-Breiman Theorem and Kingman’s subadditive ergodic theorem (applied to the sequence $\omega\mapsto g_{n}(\omega):=-\log\|(f^{\lambda}_{\omega|_{n}})^{\prime}\|$).
To show the lower bound, let us apply Theorem 3.11. That is, for $\eta$-almost every $\lambda\in U$ such that $s({\mathcal{F}}^{\lambda})<d$ we get that for every ergodic measure, almost every point has exponential distance from the enemy. Namely, for $\mu$-almost every $u$ and for every $\varepsilon>0$ there exists $C>0$ such that
$B(\Pi_{\lambda}(u),\|(f^{\lambda}_{u|_{n}})^{\prime}\|^{1+\varepsilon})\cap\bigcup_{\begin{subarray}{c}v\in{\mathcal{A}}^{n}\\ v\neq u|_{n}\end{subarray}}\Pi_{\lambda}([v])=\emptyset.$
Hence,
$\Pi_{\lambda}\nu\left(B(\Pi_{\lambda}(u),\|(f^{\lambda}_{u|_{n}})^{\prime}\|^{1+\varepsilon})\right)\leq\nu([u|_{n}]),$
and so, similarly to the upper bound,
$\underline{d}(\Pi_{\lambda}\nu,x)=\liminf_{n\to\infty}\frac{\log(\Pi_{\lambda})_{*}\nu\left(B(\Pi_{\lambda}(u),\|(f^{\lambda}_{u|_{n}})^{\prime}\|^{1+\varepsilon})\right)}{\log\|(f^{\lambda}_{u|_{n}})^{\prime}\|^{1+\varepsilon}}\geq\liminf_{n\to\infty}\frac{\log\nu\left([u|_{n}]\right)}{\log\|(f^{\lambda}_{u|_{n}})^{\prime}\|^{1+\varepsilon}}=\frac{h(\mu)-h(\mu\|\nu)}{(1+\varepsilon)\chi(\mu,\lambda)}.$
Since $\varepsilon>0$ was arbitrary, the claim follows. ∎
@@
Let =${\mathcal{F}}=\{f_{i}(x)=\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be a self-similar IFS on $\mathbb{R}^{d}$ with similarity dimension $s_{0}=s({\mathcal{F}})$, let $\underline{p}=(p_{i})_{i\in{\mathcal{A}}}$ be a probability vector and let $\mu=\underline{p}^{\mathbb{N}}$ be the Bernoulli measure corresponding to $\underline{p}$. If $(p_{i})_{i\in{\mathcal{A}}}\neq(\lambda_{i}^{s_{0}})_{i\in{\mathcal{A}}}$ and $s_{0}<d$ then for every $\alpha\in\left(\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max_{i}\frac{\log p_{i}}{\log\lambda_{i}}\right]$
$\dim_{H}\left\{x:d(\Pi\mu,x)=\alpha\right\}\leq\inf_{q\leq 0}(\alpha q+T(q))=T^{*}(\alpha),$
where $T(q)$ is the unique solution of $\sum_{i\in{\mathcal{A}}}p_{i}^{q}\lambda_{i}^{T(q)}=1$.
The proof of the lemma is standard, but for completeness, we give here a proof of this lemma.
@@
By definition,
$\left\{x:d(\Pi_{*}\mu,x)=\alpha\right\}=\bigcap_{p=1}^{\infty}\bigcup_{N=1}^{\infty}\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\Pi_{*}\mu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}.$
@@
(8.2) $\dim_{H}\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\Pi_{*}\mu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}\leq q(\alpha-1/p)+T(q)+\varepsilon.$
Let $X_{N,p}=\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\Pi_{*}\mu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}$ and $\mathcal{B}^{N}=\{B(x,2^{-n}):n\geq N\text{ and }x\in X_{N,p}\}$. Then by Besicovitch’s covering theorem (see for example [BSS23, Theorem B.3.2]), there exists $Q=Q(d)$ such that for every $i=1,\ldots,Q$ there exists a countable family ${\mathcal{B}}_{i}^{N}\subseteq{\mathcal{B}}^{N}$ such that for every $B,B^{\prime}\in{\mathcal{B}}_{i}^{N}$, $B\cap B^{\prime}=\emptyset$ and $X_{N,p}\subseteq\bigcup_{i=1}^{Q}\bigcup_{B\in{\mathcal{B}}^{N}_{i}}B$. Clearly, $X_{N,p}\subseteq X_{N+1,p}$, and so $\bigcup_{i=1}^{Q}{\mathcal{B}}_{i}^{M}$ is a cover for $X_{N,p}$ for every $M\geq N$. Hence,
$\begin{split}\mathcal{H}_{2^{-M}}^{q(\alpha-1/p)+T(q)+\varepsilon}(X_{N,p})&\leq\sum_{i=1}^{Q}\sum_{B\in{\mathcal{B}}_{i}^{M}}|B|^{q(\alpha-1/p)+T(q)+\varepsilon}\leq\sum_{i=1}^{Q}\sum_{B\in{\mathcal{B}}_{i}^{M}}\Pi_{*}\mu(B)^{q}|B|^{T(q)+\varepsilon}=\star.\end{split}$
@@
By the result of Barral and Feng [BF21], it is enough to show that (3.11) holds for $\alpha\in\left(\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max_{i}\frac{\log p_{i}}{\log\lambda_{i}}\right]$. The upper bound follows by Lemma 8.2.
If $\alpha\in\left(\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max_{i}\frac{\log p_{i}}{\log\lambda_{i}}\right)$ then there exists unique $q\leq 0$ such that $T^{\prime}(q)=-\alpha$. Moreover,
$T^{\prime}(q)=-\frac{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log p_{i}}{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log|\lambda_{i}|}\text{ and }q\alpha+T(q)=-qT^{\prime}(q)+T(q)=T^{*}(\alpha).$
see for example [Fal97, Chapter 11] or [BSS23, Chapter 5].
To show the lower bound, let us apply Lemma 8.1 with $\nu$ being the Bernoulli measure with the probabilities $(p_{i})_{i\in{\mathcal{A}}}$ and $\mu$ being the Bernoulli measure with the probabilities $(p_{i}^{q}|\lambda_{i}|^{T(q)})_{i\in{\mathcal{A}}}$. That is, let $Z\subset\mathbb{R}^{N}$ be a full measure subset such that (8.1) holds for every ergodic measure. So, for $\mu$-almost every $u$
$d(\Pi_{\underline{t}}\nu,\Pi_{\underline{t}}(u))=\frac{h(\mu)-h(\mu\|\nu)}{\chi(\mu)}=\frac{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log p_{i}}{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log|\lambda_{i}|}=\alpha.$
Hence,
$\dim_{H}\{x:d(\Pi_{\underline{t}}\nu,x)=\alpha\}\geq\dim_{H}(\Pi_{\underline{t}})_{*}\mu=\frac{h(\mu)}{\chi(\mu)}=-q\frac{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log p_{i}}{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log|\lambda_{i}|}+T(q)=q\alpha+T(q)=T^{*}(\alpha).$
@@
To complete the proof, if $\alpha=\max_{i}\frac{\log p_{i}}{\log|\lambda_{i}|}$ then let ${\mathcal{A}}_{\max}=\left\{i\in{\mathcal{A}}:\frac{\log p_{i}}{\log|\lambda_{i}|}=\alpha\right\}$ and one can repeat the argument above with $\mu$ being the uniformly distributed Bernoulli measure on ${\mathcal{A}}_{\max}^{\mathbb{N}}$. ∎
Let us note that for higher dimensional systems, we can also extend the previous result of Barral and Feng [BF21, Theorem 1.2].
Theorem 8.3.
Let ${\mathcal{A}}$ be a finite set and for each $i\in{\mathcal{A}}$ fix $\lambda_{i}\in(0,1)$ and a $d\times d$ orthogonal matrix $O_{i}$. For $t=(t_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{{\mathcal{A}}}$ let ${\mathcal{F}}_{t}=\{f_{i}(x)=\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be a self-similar IFS on the line. Let $s_{0}=s({\mathcal{F}}_{t})$ be the similarity dimension of ${\mathcal{F}}_{t}$ and assume that $s_{0}<d$ and $\max_{i}\lambda_{i}<1/2$ Then the following holds for Lebesgue almost every $(t_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{\mathcal{A}}$. For every probability vector $\underline{p}=(p_{i})_{i\in{\mathcal{A}}}$ such that $(p_{i})_{i\in{\mathcal{A}}}\neq(\lambda_{i}^{s_{0}})_{i\in{\mathcal{A}}}$, the multifractal formalism (3.11) holds for the self-similar measure $\Pi_{{\mathcal{F}}_{t}}(\underline{p}^{\mathbb{N}})$ and every $\alpha\in\left[\frac{\sum_{i}p_{i}\log p_{i}}{\sum_{i}p_{i}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$ simultaneously.
The theorem follows by the combination of [BF21, Theorem 1.2(2)], Lemma 8.1, Lemma 8.2 and Example 3.8, and left for the reader.

### v3

@@ 7.2. Gibbs measures
where $\|\cdot\|$ denotes the supremum norm on $\Sigma$ (see also [Wal82, Theorem 9.7.(iv)]). Combining this with (7.8) gives that for every $\phi_{1},\phi_{2}$ there exists $M=M(\phi_{1},\phi_{2})$ such that for every $\omega\in\Sigma$ and $n\in{\mathbb{N}}$
@@
$\omega|_{n+1}]\subset B(\omega,r)\subset[\omega|_{n}]\text{ and }B(\omega,2r)\subset[\omega|_{n-N}].$
@@ 8. Multifractal spectrum of self-similar measures
In this section we prove Theorem 3.15. Let us begin with a general consequence of the nearly bi-Lipschitz property on the local dimensions of projections. Recall that for conformal iterated function systems, the nearly bi-Lipschitz property is equivalent to the EDE separation condition (Proposition 3.10).
@@
Let $(X,\rho_{X})$ and $(Y,d_{Y})$ be metric spaces. Let $\Pi:X\to Y$ be a Lipschtiz map and let $\mu$ be a finite Borel measure on $X$. If $\Pi$ is $\mu$-nearly bi-Lipschitz, then the following holds for $\mu$-a.e $x\in X$:
[…]
@@
As we assume that $\Pi$ is Lipschitz, inequalities
[…]
hold for every $x\in X$ and every finite Borel measure $\nu$ on $X$. It therefore remains to prove the opposite inequalities for $\mu$-a.e. $x\in X$. Let $x\in X$ and $\alpha\in(0,1)$ be such that there exists $C=C(x,\alpha)$ so that
(8.2) $\rho_{X}(x,y)\leq C\rho_{Y}(\Pi(x),\Pi(y))^{\alpha}\text{ for every }y\in X.$
Then for all $r>0$
$\Pi^{-1}(B(\Pi(x),r))\subset B(x,Cr^{\alpha}),$
so for any finite Borel measure $\nu$ on $X$
$\Pi\nu(B(\Pi(x),r))\leq\nu(B(x,Cr^{\alpha})).$
Therefore, for $0<r<1$
$\frac{\log\Pi\nu(B(\Pi(x),r))}{\log r}\geq\frac{\log\nu(B(x,Cr^{\alpha}))}{\log r}=\frac{\log Cr^{\alpha}}{\log r}\cdot\frac{\log\nu(B(x,Cr^{\alpha}))}{\log Cr^{\alpha}}.$
Taking $\liminf\limits_{r\to 0}$ and $\limsup\limits_{r\to 0}$ yields
[…]
If $\Pi$ is $\mu$-nearly bi-Lipschitz, then for $\mu$-a.e. $x\in X$, inequality (8.2) holds for every $\alpha\in(0,1)$, and hence (8.3) gives
[…]
for $\mu$-a.e. $x$ and every $\nu$. Combined with (8.1), this finishes the proof. ∎
Now we can turn to proving Theorem 3.15. We will use the following notation: for a self-similar measure $\nu$ on $\mathbb{R}^{d}$ corresponding to the IFS ${\mathcal{F}}=\{x\mapsto\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ with $\lambda_{i}\in(0,1)$, $d\times d$ orthogonal matrices $O_{i}$ and $t_{i}\in\mathbb{R}$ and a probability vector $p=(p_{i})_{i\in{\mathcal{A}}}$, given $\alpha\in\mathbb{R}$ we set
$T^{*}(\alpha):=\inf_{q\in\mathbb{R}}(\alpha q+T(q)),$
where $T(q)$ is the unique real solution of $\sum_{i\in{\mathcal{A}}}p_{i}^{q}\lambda_{i}^{T(q)}=1$. First, we establish the lower bound.
@@
Let ${\mathcal{F}}=\{f_{i}(x)=\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be an IFS consisting of similarities of $\mathbb{R}^{d}$, with similarity dimension $s_{0}=s({\mathcal{F}})$, let $p=(p_{i})_{i\in{\mathcal{A}}}$ be a probability vector and let $\nu$ be the corresponding self-similar measure. If $(p_{i})_{i\in{\mathcal{A}}}\neq(\lambda_{i}^{s_{0}})_{i\in{\mathcal{A}}}$ and $s_{0}<d$ then for every $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$
$\dim_{H}\left(\left\{x:d(\nu,x)=\alpha\right\}\right)\leq T^{*}(\alpha)=\inf_{q\leq 0}(\alpha q+T(q))$
The proof of the lemma is standard, but we include it for completeness.
@@
If $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right)$, then there exists unique $q=q_{\alpha}\leq 0$ such that
(8.4) $T^{\prime}(q)=-\alpha,$
and moreover
(8.5) $T^{\prime}(q)=-\frac{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log p_{i}}{\sum_{i}p_{i}^{q}|\lambda_{i}|^{T(q)}\log|\lambda_{i}|}\text{ and }q\alpha+T(q)=-qT^{\prime}(q)+T(q)=T^{*}(\alpha),$
see for example [Fal97, Chapter 11] or [BSS23, Chapter 5]. Consequently
(8.6) $T^{*}(\alpha)=\inf_{q\leq 0}(\alpha q+T(q)).$
By continuity, (8.6) extends to $\alpha=\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}$. It therefore suffices to show that for every $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$ and $q\leq 0$ inequality
$\dim_{H}\left(\left\{x:d(\nu,x)=\alpha\right\}\right)\leq\alpha q+T(q)$
holds.
Let $\mu=p^{\mathbb{N}}$ be the Bernoulli measure corresponding to $p$. By definition,
$\left\{x:d(\nu,x)=\alpha\right\}=\bigcap_{p=1}^{\infty}\bigcup_{N=1}^{\infty}\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\nu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}.$
@@
(8.7) $\dim_{H}\left(\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\nu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}\right)\leq q(\alpha-1/p)+T(q)+\varepsilon.$
Let $X_{N,p}=\bigcap_{n=N}^{\infty}\{x:2^{-n(\alpha+1/p)}\leq\nu(B(x,2^{-n}))\leq 2^{-n(\alpha-1/p)}\}$ and $\mathcal{B}^{N}=\{B(x,2^{-n}):n\geq N\text{ and }x\in X_{N,p}\}$. Then by Besicovitch’s covering theorem (see for example [BSS23, Theorem B.3.2]), there exists $Q=Q(d)$ such that for every $i=1,\ldots,Q$ there exists a countable family ${\mathcal{B}}_{i}^{N}\subseteq{\mathcal{B}}^{N}$ such that for every distinct $^{\prime}\in{\mathcal{B}}_{i}^{N}, BB=B\cap B^{\prime}$$\emptyset$ and $X_{N,p}\subseteq\bigcup_{i=1}^{Q}\bigcup_{B\in{\mathcal{B}}^{N}_{i}}B$. Clearly, $X_{N,p}\subseteq X_{N+1,p}$, and so $\bigcup_{i=1}^{Q}{\mathcal{B}}_{i}^{M}$ is a cover for $X_{N,p}$ for every $M\geq N$. Hence,
$\begin{split}\mathcal{H}_{2^{-M}}^{q(\alpha-1/p)+T(q)+\varepsilon}(X_{N,p})&\leq\sum_{i=1}^{Q}\sum_{B\in{\mathcal{B}}_{i}^{M}}|B|^{q(\alpha-1/p)+T(q)+\varepsilon}\leq\sum_{i=1}^{Q}\sum_{B\in{\mathcal{B}}_{i}^{M}}\nu(B)^{q}|B|^{T(q)+\varepsilon}=\star.\end{split}$
@@
Our goal is to prove that for almost every $t$, equality
(8.8) $\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)=\inf_{q\in\mathbb{R}}(\alpha q+T(q))$
holds simultaneously for all $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$. This will finish the proof in the case $d>1$. For $d=1$ we can additionally apply [BF21, Theorem 1.2, Remark 7.3], showing that for almost every $t$, (8.8) holds for $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}}\right]$. The upper bound $\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)\leq\inf\limits_{q\in\mathbb{R}}(\alpha q+T(q))$ follows by Lemma 8.2.
We shall prove the lower bound for all $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$. Let us first introduce some notation. Let $\Sigma={\mathcal{A}}^{N}$ be the symbolic space. Given $\alpha\in\left(\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right)$, let $q_{\alpha}\in\mathbb{R}$ be such that (8.4) and (8.5) hold with $q=q_{\alpha}$ (recall e.g. [BSS23, Chapter 5]). Further, for a given probability vector $p=(p_{i})_{i\in{\mathcal{A}}}$, let $\nu^{p}=p^{\otimes{\mathbb{N}}}$ be the Bernoulli measure on $\Sigma$ corresponding to $p$ and let $\mu_{\alpha,p}$ be the Bernoulli measure on $\Sigma$ corresponding to the probability vector $(p_{i}^{q_{\alpha}}|\lambda_{i}|^{T(q_{\alpha})})_{i\in{\mathcal{A}}}$. Let $\Pi_{t}:\Sigma\to\mathbb{R}^{d}$ be the natural projection map corresponding to the IFS ${\mathcal{F}}_{t}$. Note that with this notation $\nu_{t,p}=\Pi_{t}\nu^{p}$. Endow $\Sigma$ with the adapted metric $\rho=\rho_{{\mathcal{F}}_{t}}$ defined in (3.3) and corresponding to the IFS ${\mathcal{F}}_{t}$ (note that $\rho_{{\mathcal{F}}_{t}}$ does not depend on $t$; in fact it depends only on contractions $\lambda_{i},i\in{\mathcal{A}}$, which we treat as fixed). A direct computation (see e.g. [BSS23, Lemma 5.1.3]) shows that (with respect to metric $\rho$ on $\Sigma$)
(8.9) $d(\nu^{p},\omega)=\alpha\text{ for }\mu_{\alpha,p}\text{-a.e. }\omega\in\Sigma$
and, by (3.4), (8.4) and (8.5)
(8.10) $\dim_{H}\mu_{\alpha,p}=\frac{h(\mu_{\alpha,p})}{\chi(\mu_{\alpha,p})}=-q_{\alpha}\frac{\sum_{i}p_{i}^{q_{\alpha}}|\lambda_{i}|^{T(q_{\alpha})}\log p_{i}}{\sum_{i}p_{i}^{q_{\alpha}}|\lambda_{i}|^{T(q_{\alpha})}\log|\lambda_{i}|}+T(q_{\alpha})=q_{\alpha}\alpha+T(q_{\alpha})=T^{*}(\alpha).$
By Example 3.8, the assumptions of Theorem 3.15 are sufficient to establish the transversality condition for the family ${\mathcal{F}}_{t}$ with $t$ as the parameter. We can therefore apply Theorem 2.7 (or, more directly, Theorem 3.12) and, as we assume $s_{0}=s({\mathcal{F}}_{t})<d$, we can conclude that for $\eta$-a.e. $t\in(\mathbb{R}^{d})^{{\mathcal{A}}}$, simultaneously for every $\alpha$ and $p$, the natural projection $\Pi_{t}$ is $\mu_{\alpha,p}$-nearly bi-Lipschitz, and (recalling (8.10)) equality
(8.11) $\dim_{H}\Pi_{t}\mu_{\alpha,p}=T^{*}(\alpha)$
holds. It therefore follows from Lemma 8.1 and (8.9) that for $\eta$-a.e. $\lambda$ and every $\alpha,p$
$d(\nu_{t,p},\Pi_{t}(\omega))=d(\Pi_{t}\nu^{p},\Pi_{t}(\omega))=d(\nu^{p},\omega)=\alpha\text{ for }\mu_{\alpha,p}\text{-a.e.}\omega\in\Sigma.$
This implies (as the above shows $\Pi_{t}\mu_{\alpha,p}\left(\{x:d(\nu_{t,p},x)=\alpha\}\right)=1$)
$\dim_{H}\left(\{x:d(\nu_{t,p},x)=\alpha\}\right)\geq\dim_{H}\Pi_{t}\mu_{\alpha,p}$
and so by (8.11)
(8.12) $\dim_{H}\left(\{x:d(\nu_{t,p},x)=\alpha\}\right)\geq T^{*}(\alpha)$
This establishes (8.8) for $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right)$.
@@
To complete the proof, if $\alpha=\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}$ then in order to obtain the lower bound (the upper bound in this case follows from Lemma 8.2), one can let ${\mathcal{A}}_{\max}=\left\{i\in{\mathcal{A}}:\frac{\log p_{i}}{\log|\lambda_{i}|}=\alpha\right\}$ and repeat the argument above with $\mu_{\alpha,p}$ being the uniformly distributed Bernoulli measure on ${\mathcal{A}}_{\max}^{\mathbb{N}}$. This establishes (8.8) for $\alpha\in\left[\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$. For $\alpha=\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}$, one can similarly repeat the argument to obtain the lower bound (8.12).
Note that it was crucial for the above proof that for almost every translation parameter $t$, the $\mu_{\alpha,p}$-nearly bi-Lipschitz property (i.e. EDE) holds simultaneously for all $\alpha$ (and $p$), as follows from the universal projection theorem (Theorem 2.7). If it was established only for fixed $\alpha$, one would obtain equality (8.8) only for a single level-set rather then on the full spectrum $\left(\frac{\sum_{i}\lambda_{i}^{s_{0}}\log p_{i}}{\sum_{i}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$.